• Suppose FCFF = $1 mil for years 1-4 and then is expected to grow at a rate of 3%. Assume WACC = 15%. If 500,000 shares are outstanding, what is the predicted price of this stock if the firm has $5,000,000 of debt?
Answer #1

Year 5 FCFF = 1,000,000 (1 + 3%) = 1,030,000

Value at year 4 = FCFF5 / required rate - growth rate

Value at year 4 = 1,030,000 / 0.15 - 0.03

Value at year 4 = 1,030,000 / 0.12

Value at year 4 = 8,583,333.333

Total firm value = Annuity * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Total firm value = 1,000,000 * [1 - 1 / (1 + 0.12)4] / 0.12 + 8,583,333.333 / (1 + 0.12)4

Total firm value = 1,000,000 * [1 - 0.635518] / 0.12 + 5,454,863.506

Total firm value = 1,000,000 * 3.037349 + 5,454,863.506

Total firm value = $8,492,212.853

Equity value = 8,492,212.853 - 5,000,000

Equity value = 3,492,212.853

Price of stock = 3,492,212.853 / 500,000

Price of stock = $6.98